Accelerator Update for February 27, 2008 - February 29, 2008
Wednesday February 27
The day shift began with Operations monitoring store 5934 (64.44E30), the Recycler maintaining a stash (290.3E10), with the Antiproton Source (Pbar) stack at (11.6E10), with the Linac, Booster, Main Injector (MI), and BooNEs off because of problems with LRF3.
LRF3 was repaired by 2:44 PM.
Beams returns to Pbar and BooNEs at 3:00 PM.
Operations terminated store 5934 at 5:00 PM.
Tevatron Quench at B3 occurred at 6:47 PM.
At 10:19 PM the Tev was ramping once again.
Operations put the Tevatron (TeV) into shot setup at 11:04 PM.
Thursday February 28
The midnight shift began with the Tevatron in shot setup and with BooNEs taking beam.
At 12:36 AM, the Tevatron quenched at E39. Recovery and investigation were underway.
Linac trim magnets tripped off at 2:50 AM, experts were called.
Linac trim magnets were repaired at 6:00 AM.
Stacking and beam to the BooNEs resumed at 6:21 AM.
Controls problems with the Tevatron A Front End began occurring at 8:54 AM.
These problems hampered Tevatron tuning throughout the dayshift.
Linac KRF5 tripped at 4:14 PM, experts investigated.
A bad solenoid power supply was changed for KRF5 and beam returns to Pbar and the BooNEs at 5:24 PM.
Operations put the Tevatron (TeV) into shot setup at 7:26 PM.
Operations established store 5943 at 11:38 PM with an initial luminosity of 180.7 E30.
Friday February 29
The midnight shift began with Operations monitoring store 5943 (120.86E30), the Recycler maintaining a stash (105.6E10), with the Antiproton Source (Pbar) stack at (39.1E10), and beam to the BooNEs.
The Plans for Friday and the weekend
Continue to stack, stash, and store.
Comments and Suggestions
|last modified 02/29/2008 email Fermilab|
|Security, Privacy, Legal|